# Diamond-loaded polyimide aerogel scattering filters and their applications in astrophysical and planetary science observations

### v1

@@ II Comparison of Infrared-Blocking Filter Technologies
Figure 2: Top: Comparison of the attenuation coefficient of a variety of infrared blocking filter materials. The Zotefoam (ZF) samples, XPS foam, and unloaded aerogels all have comparable attenuation. The data for the Zitex sample below 4 THz are binned because the thinness of the sample produces significant etalon fringing in the measurement data. Bottom: Comparison of the transmittance of different infrared filters. The Zotefoam filter and XPS foams are shown for a filter 3 mm thick. The Zitex is shown for a filter that is 0.1 mm thick. The diamond-loaded polyimide aerogel filter is loaded with 60 mg/cc of 10-20 µm and 40 mg/cc of 3-6 µm particles, and is a sample that is 0.18 mm thick. The metal-mesh filter is 0.004 µm thick.
@@
Microporous PTFE, such as ZitexSaint-Gobain Tape Solutions (2025) and Porex, [^1] Porex Filtration Group, https://www.porex.com/products/ is also a promising candidate material for use as an IR-blocking filter. A range of formulations are commercially available with typical pore sizes down to 1 µm, producing cutoff frequencies due to scattering from the pore structure to shorter wavelengths than for expanded foams; however, the intrinsic absorption of PTFE at THz frequencies limits use of these materials to the microwave and sub-millimeter. Similar to expanded polymer foams, the properties of microporous PTFE are not particularly tunable to the desired wavelength range of the end user. It can be procured in a variety of thicknesses, but the cutoff frequency is largely set by the pore size, high index of refraction (n $\simeq 1.44$), and high loss of PTFE in the THz.
@@
One specific metal-mesh filter architecture is designed to reject infrared light near the peak emission of a 300 K (or other specified temperature) blackbody. These filters are called “infrared shaders” Tucker and Ade (2006) and are constructed of a single thin, $\sim 3$ µm polypropylene film with capacitive grids on both sides. These grids can be seen in microscope images in Figure 4.
@@ III.1 Total Hemispherical Measurements of Diamond-loaded Polyimide Aerogel Filters and Other Media
Figure 6: Comparison of the measured total hemispherical transmittance (top panel), measured total hemispherical reflectance (middle panel), and estimated emissivity (bottom panel) of different infrared filters. The Zotefoam filter and XPS foams are shown for a filter 3 mm thick. The Zitex is shown for a filter that is 0.1 mm thick. The diamond-loaded polyimide aerogel filter is an EXCLAIM candidate formulation, Formula G from 2. The sample is loaded with 60 mg/cc of 10-20 µm and 40 mg/cc of 3-6 µm particles and is 0.18 mm thick. The metal-mesh filter is 3.3 µm thick. The white LD24 and HD30 samples show higher transmission than the other samples, which leads to a lower calculated emissivity.
@@ IV The Fabrication of large-scale Aerogel Filters
Polyimide aerogel films were prepared from biphenyl-3,3’,4,4’-tetracarboxylic dianhydride (BPDA) and combination of 2,2’-dimethylbenzidine (DMBZ, 75 mol%) and 4,4’-Bis (4-aminophenoxy) biphenyl (BAPB, 25 mol%) then crosslinked by cross-linker 3,5-triaminophenoxybenzene (TAB) in N-methyl-2-pyrrolidinone (NMP). The polymer concentration is formulated to 10% by weight. The molar ratio of BPDA and the total diamines is 26:25. For example, to make a sample film 460 mm x 460 mm with 1 mm thickness, 5.94 g BAPB, and 19.61 g BPDA were mixed and stirred in 326.03 mL NMP for 30 minutes. 10.27 g DMBZ in 20 ml NMP was added in the mixture until the sol cleared and followed by addition of 0.57 g TAB in 10 mL NMP. Diamond particles with different sizes and loading amounts were added in this stage prior to adding 50.42 mL acetic anhydride and then 4.65 mL triethylamine. The polyimide gel films undergo supercritical fluid extraction using liquid CO2 to obtain the final aerogel films.
We investigaed three different methods for fabricating large area films. The first was a vertical casting mold, which is constructed of two rigid pieces of plastic separated by spacers of a prescribed thickness. After stirring for 15 minutes, the sol was poured into a vertical mold 52.5 cm x 52.5 cm with the spacer thickness 1.2 mm, thus the opening is 52.5 cm x 1.2 mm.
@@ V Designing filters for specific mission requirements
Aerogel filter samples were also tested for their cryogenic durability and survivability. Many experiments and instruments observing the THz through millimeter-wave regimes use cryogenic detectors requiring optics cooled to below 77 K, and in some cases below 4 K. Samples were repeatedly immersed in liquid nitrogen to test thermal shock performance. Filter samples were also installed inside of a cryogenic testbed and repeatedly cycled down to 4 K from room temperature, as described in Section VI. While the filter mount must be designed to accommodate the CTE-mismatch between the polymer-based aerogels and other metallic or semi-metallic components, our polyimide aerogel filters are robust against thermal shock and cryogenic cycling.
@@ V.1 Aerogel Filter Design Principles
The diamond particle size and loading density are varied in order to tune the filter to the desired cutoff frequency. Particle size has the strongest effect on cutoff frequency. Larger particles have a lower cut-off frequency and smaller particles have a higher cut-off frequency. Diamond powders used typically have a Gaussian size distribution with the named range, e.g. 3-6 µm, representing the 3-$\sigma$ size range. Polyimide aerogel filters have been fabricated with a range of diamond loading densities ranging from 20 mg/cc up to 100 mg/cc. Loading densities significantly above 100 mg/cc prevent proper polymerization and curing of the base aerogel. Figure 8 shows a summary of the parameter space generally available to us when designing and manufacturing diamond-loaded filters. For a given particle size distribution, a higher loading density creates a filter with a lower cut-off frequency and vice versa. In general, the scattering cross-section of the diamond particles goes as the square of the particle radius (area $\propto r^{2}$) and mass density goes as the cube of particle radius (volume $\propto r^{3}$). In order to capture the effect of particle radius and loading density, a column density can be calculated, that represents the total area of scatterers per unit area. Multiple particle size distributions and densities can be mixed together, e.g. primarily using a higher density of larger particles to set the primary cut-off frequency and mixing in lower densities of smaller particles to control high frequency leakage.
@@ V.2 Terahertz Applications
Figure 9: Transmission data for five different diamond-loaded polyimide aerogel filters for the SSOLVE mission. Filter loading schema are in Table 2. Because not all of the prototype filters are the same thickness, all filter transmission data were scaled to represent a 200 µm thick sample. Black vertical dashed line is at 2.5 THz, and the red circle represents 90% transmission.
@@ V.3 Microwave and Sub-Millimeter Applications
Figure 10 shows several different prototype filters for these two instruments. All of these filters cut off most of the infrared light above 4 THz while yielding nearly unity transmission below 1 THz. Formula G, likely due to the addition of smaller particles, shows lower transmission above 10 THz than the other two formulas.
@@ VI Thermal modeling of filters in a receiver
To predict the thermal performance of the aerogel filters in a cryogenic receiver, we simulate radiative heat transfer and heat conduction of the filter stack using finite element analysis with COMSOL Multiphysics. Figure 11 shows the 2D geometry of the model and the simulated equilibrium temperature under the operating conditions of a CLASS cryogenic receiver. The final 3D geometry is generated by revolving the 2D geometry about the axis of revolution, which is denoted by the blue dashed line. The CLASS cryogenic receiver employs a multistage design, where each stage is referred to by its nominal temperature, namely the 300 K stage, 60 K stage, and 4 K stage. The focal plane (the coldest part of the cryogenic receiver) is located between the bottom 4 K aerogel filter and the 4K baseplate. The two different filter configurations considered here are:
@@
Multiple configurations were tested because the exact emissivity and thus the exact thermal load in each scenario was unknown prior to the measurements. Both configurations include five foam filters on the 300 K stage, denoted by he white lines under the window in Figure 11. Note that, for simplification, the modeled 60 K stage and 300 K stage are connected as one geometry instead of two, and the length of the modeled receiver is also shorter than that of the actual receiver.
@@
These simulations consider both thermal conduction and radiation inside the region of interest. The aerogel filters are treated as uniform, perfect diffusers that they transmit, radiate, and reflect uniformly in all directions. Simulations are performed iteratively for different emissivity values ranging from 0 to 0.3. Figure 12 shows the temperature and the total emitted power of the bottom aerogel filter at the 4 K stage as a function of emissivity. The blue line and the orange line represent the result in a three-filter configuration and five-filter configuration respectively. The red dashed horizontal line indicates the upper limit of load (0.1 W) required by the CLASS experiment.
@@
Filter configurations Filter Temperature (K)
4 K Baseplate
Temperature (K)
4 K Baseplate
Load (W)
@@ VII Cryogenic performance of Aerogel filters
where $T$ is the temperature in Kelvin and $P$ is the power in watts. The 4 K baseplate load inferred from the above load curve for each filter configuration is summarised in the rightmost column in Table 3. Note that this inferred power is much higher than the power emitted by the bottom 4 K filter as simulated by COMSOL. This could be because the thermal load from the window reached the baseplates through other paths, such as the gap between shields of adjacent stages, instead of going through the aerogel filter only. Another possible reason is that there are tiny air bubbles present in the aerogel filters that make them slightly more transmissive in bulk than optical measurements suggest, allowing more load to pass onto the baseplates than expected.

### v2

@@ II Comparison of Infrared-Blocking Filter Technologies
Figure 2: Top: Comparison of the attenuation coefficient of a variety of infrared blocking filter materials. The Zotefoam (ZF) samples, XPS foam, and unloaded aerogels all have comparable attenuation. The data for the Zitex sample below 4 THz are binned because the thinness of the sample produces significant etalon fringing in the measurement data. Bottom: Comparison of the transmittance of different infrared filters. The Zotefoam filter and XPS foams are shown for a filter 3 mm thick. The Zitex is shown for a filter that is 0.1 mm thick. The metal-mesh filter is 0.004 µm thick. The low frequency aerogel is Formula G and the high frequency aerogel is Formula C from Table 2.
@@
Microporous PTFE, such as ZitexSaint-Gobain Tape Solutions (2025) and Porex, [^1] Porex Filtration Group, https://www.porex.com/products/ is also a promising candidate material for use as an IR-blocking filter. A range of formulations are commercially available with typical pore sizes down to 1 µm. Scattering from the small micropore structure produces filter cutoffs at higher frequencies/shorter wavelengths than for expanded foams; however, the intrinsic absorption of PTFE at THz frequencies limits use of these materials to the microwave and sub-millimeter. Similar to expanded polymer foams, the properties of microporous PTFE are not particularly tunable to the desired wavelength range of the end user. It can be procured in a variety of thicknesses, but the cutoff frequency is largely set by the pore size, high index of refraction (n $\simeq 1.44$), and high loss of PTFE in the THz.
@@
One specific metal-mesh filter architecture is designed to reject infrared light near the peak emission of a 300 K (or other specified temperature) blackbody. These filters are called “infrared shaders” Tucker and Ade (2006) and are constructed of a single thin, $\sim 3$ µm polypropylene film with capacitive grids on both sides. These grids can be seen in microscope images in Figure 4. They have feature sizes that are approximately 20 µm wide with spacing that is less than 10 µm.
@@
Scattering filters like foams and aerogels reject unwanted light somewhat differently than metal-mesh filters. Metal-mesh filters are mostly specular and reflect light backwards through an optical system. This reflected light must be able to pass backwards through the preceding optical elements. Foams and scattering filters will reflect light more isotropically into the surrounding environment. These considerations must be taken into consideration when desiging an instrument around a certain filter type.
@@ III.1 Total Hemispherical Measurements of Diamond-loaded Polyimide Aerogel Filters and Other Media
Figure 6: Comparison of the measured total hemispherical transmittance (top panel), measured total hemispherical reflectance (middle panel), and estimated emissivity (bottom panel) of different infrared filters. The Zotefoam filter and XPS foams are shown for a filter 3 mm thick. The Zitex is shown for a filter that is 0.1 mm thick. The diamond-loaded polyimide aerogel filter is an EXCLAIM candidate formulation, Formula G from Table 2, and is 1.3 mm thick. The unloaded aerogel is 1.4 mm thick. The metal-mesh filter is an infrared shader 3.3 µm thick. The white LD24 and HD30 samples show higher transmission than the other samples, which leads to a lower calculated emissivity.
@@ IV The Fabrication of large-scale Aerogel Filters
The most widely practiced procedure in polyimide synthesis is the two-step polymerization method which involves an intermediate step forming a poly(amic acid) along the reaction pathway to forming the polyimide. Creating a polymer aerogel using this method requires the use of sol-gel polymerization by combining bi-functional monomers and solvent to form a colloidal solution forming a dilute polymer network within the solution, leading to gelation. The gel networks created by sol-gel method become the aerogel skeletal structure once the liquid portion is removed, creating the aerogel.
To create a polyimide aerogel, equimolar amounts of dianhydrides and diamines, in a polar aprotic solvent (n-methyl-2-pyrrolidone (NMP)) first form a poly(amic) acid precursor, followed by chemical imidization utilizing a base catalyst (triethyl amine) and a water scavenger (acetic anhydride) and then cross-linked with a multifunctional monomer.
The diamines, 2,2’-dimethylbenzidine (DMBZ) and 4,4’-Bis (4-aminophenoxy) biphenyl (BAPB), are first added to NMP to create a colloidal solution (sol) of suspended monomers in solvent. To this solution the dianhydride, biphenyl-3,3’,4,4’-tetracarboxylic dianhydride (BPDA), is added in excess to react with the diamine to create the soluble amic acid oligomers with anhydride endcaps. The triamine, 3,5-triaminophenoxybenzene (TAB), is added to create a crosslinked polymer network followed by chemical imidization to initiate ring closure of the amic acid to form a crosslinked polyimide gel. The polyimide gel films undergo supercritical fluid extraction using liquid CO2 to remove the solvent and obtain the final rigid aerogel films.
To make a large sample film that is 460 mm x 460 mm square with 1 mm thickness, 5.94 g BAPB and 19.61 g BPDA were mixed together and stirred into 326.03 mL NMP. The whole mixture was stirred for 30 minutes. Then 10.27 g DMBZ, mixed in 20 ml NMP was added in the mixture until the sol cleared. Subsequently, 0.57 g TAB mixed in 10 mL NMP was added to the sol. At this point in the procedure, diamond particles with different sizes and loading amounts were added. After the diamond addition, 50.42 mL of acetic anhydride and 4.65 mL of triethylamine were added. Following this sol-gel synthesis, the diamond-loaded film underwent supercritical CO2 fluid extraction to produce the aerogel.
We investigated three different methods for fabricating large area films. The first was a vertical casting mold, which is constructed of two rigid pieces of plastic separated by spacers of a prescribed thickness. After stirring for 15 minutes, the sol was poured into a vertical mold 52.5 cm x 52.5 cm with the spacer thickness 1.2 mm, thus the opening is 52.5 cm x 1.2 mm. Gravity pulls the diamond particles along the gravity vector during curing, however, because of the long vertical dimension of the mold, the dispersion of large and small particles is not as substantial as it is with horizontal casting methods. The inhomogeneous sections of the top and bottom edges can be cut away and the middle homogeneous section can be used for the final product.
@@
Film homogeneity can be judged by looking at the color of the film after it is dried. Areas with substantially fewer particles will be lighter yellow; areas with more particles will be darker yellow or even a greenish-yellow. The inhomogeneities are typically caused by the largest particles, especially with sizes greater than 100 µm. The particles 3-6 µm, 10-20 µm, and 40-60 µm have fewer problems, but the 250-350 µmcause the most inhomogeneities. Five different samples were taken from the center and the four corners of a large, $\sim 50$ cm, film, with the corner samples being roughly 50 mm away from the edge of the film. The samples were characterized on the FTS and found to be uniform within the error of the measurements. Removing a 50 mm border from a 50 cm film results in an approximately 60% yield.
Our current method of drying is rolling the filters and then placing them in a supercritical dryer. The length of the film can be longer, but diameter will then be determined by the depth of the dryer chamber, the chamber at Glenn Research Center (GRC) can accommodate and approximately 55 cm sample for reasonable drying. It can be scaled up to 1 m or larger with a larger chamber.
The desired thickness of a film or filter is defined by the target application or frequency range. We made 0.2 mm and 1 mm for the above applications. Generally, the thinnest filter prepared by casting method at GRC is about 0.1 mm to make it easy to peel off from the PET film. Very thick filters have other considerations. Depending on the casting method, thicker may result in more inhomogenous filters as the differently sized particles sink at different rates through the larger depth of the film. We also developed layer-by-layer casting to get laminated films, which could be used to get very thick films with a more homogeneous distribution of particles, or a specific controlled distribution of particle sizes through the thickness of the film.
For large diameter and large thickness filters, they are also more difficult to roll to be put inside the supercritical dryer. Depending on the formulation, some samples about 4-5 mm thick can still flex about $180^{\circ}$ without breaking.
@@ V Designing filters for specific mission requirements
Aerogel filter samples were also tested for their cryogenic durability and survivability. Many experiments and instruments observing the THz through millimeter-wave regimes use cryogenic detectors requiring optics cooled to below 77 K, and in some cases below 4 K. Samples were repeatedly immersed in liquid nitrogen to test thermal shock performance. Filter samples were also installed inside of a cryogenic testbed and repeatedly cycled down to 4 K from room temperature, as described in Section VI. While the filter mount must be designed to accommodate the CTE-mismatch between the polymer-based aerogels and other metallic or semi-metallic components, our polyimide aerogel filters are robust against thermal shock and cryogenic cycling. Because the primary component of the aerogel filters is a polymer plastic, polyimide, the filters are mounted in rings made of Delrin plastic. Delrin was chosen because it is known to work well at cryogenic temperatures. These plastic rings are screwed into aluminum filter mounts, so the CTE mismatch between the aluminum and the plastic is taken up by over-sized clearance holes on the plastic filter rings, not the aerogel filters themselves.
@@ V.1 Aerogel Filter Design Principles
The diamond particle size and loading density are varied in order to tune the filter to the desired cutoff frequency. Particle size has the strongest effect on cutoff frequency. Larger particles have a lower cut-off frequency and smaller particles have a higher cut-off frequency. Diamond powders used typically have a Gaussian size distribution with the named range, e.g. 3-6 µm, representing the 3-$\sigma$ size range. Polyimide aerogel filters have been fabricated with a range of diamond loading densities ranging from 20 mg/cc up to 100 mg/cc. Loading densities significantly above 100 mg/cc prevent proper polymerization and curing of the base aerogel. The base aerogels have a density of approximately 130 mg/cc. Figure 8 shows a summary of the parameter space generally available to us when designing and manufacturing diamond-loaded filters. For a given particle size distribution, a higher loading density creates a filter with a lower cut-off frequency and vice versa. In general, the scattering cross-section of the diamond particles goes as the square of the particle radius (area $\propto r^{2}$) and mass density goes as the cube of particle radius (volume $\propto r^{3}$). In order to capture the effect of particle radius and loading density, a column density can be calculated, that represents the total area of scatterers per unit area. Multiple particle size distributions and densities can be mixed together, e.g. primarily using a higher density of larger particles to set the primary cut-off frequency and mixing in lower densities of smaller particles to control high frequency leakage.
@@ V.2 Terahertz Applications
Figure 9: Transmission data for five different diamond-loaded polyimide aerogel filters for the SSOLVE mission. Filter loading schema are in Table 2. Note that measurement data for Formulas A, B, and C only extends to approximately 20 THz, while data for Formulas D and E extend out to 1000 THz. Because not all of the prototype filters are the same thickness, all filter transmission data were scaled to represent a 200 µm thick sample. Black vertical dashed line is at 2.5 THz, and the red circle represents 90% transmission.
@@ V.3 Microwave and Sub-Millimeter Applications
Figure 10 shows several different prototype filters for these two instruments. All of these filters cut off most of the infrared light above 4 THz while yielding nearly unity transmission below 1 THz. Formula G, likely due to the addition of 10-20 µm particles, shows lower transmission above 10 THz than the other two formulas.
@@ VI Thermal modeling of filters in a receiver
To predict the thermal performance of the aerogel filters in a cryogenic receiver, we simulate radiative heat transfer and heat conduction of the filter stack using finite element analysis with COMSOL Multiphysics. Because foams and scattering filters reflect light isotropically, they are more strongly thermally-coupled to the sidewalls of a cryogenic instrument than metal-mesh filters. The stronger thermal coupling can be advantageous when combined with appropriate absorptive cryogenic coatings.
Figure 11 shows the 2D geometry of the model and the simulated equilibrium temperature under the operating conditions of a CLASS cryogenic receiver. The final 3D geometry is generated by revolving the 2D geometry about the axis of revolution, which is denoted by the blue dashed line. The CLASS cryogenic receiver employs a multistage design, where each stage is referred to by its nominal temperature, namely the 300 K stage, 60 K stage, and 4 K stage. The focal plane (the coldest part of the cryogenic receiver) is located between the bottom 4 K aerogel filter and the 4K baseplate. The two different filter configurations considered here are:
@@
Multiple configurations were tested because the exact emissivity and thus the exact thermal load in each scenario was unknown prior to the measurements. Both configurations include five foam filters on the 300 K stage, denoted by the white lines under the window in Figure 11. Note that, for simplification, the modeled 60 K stage and 300 K stage are connected as one geometry instead of two, and the length of the modeled receiver is also shorter than that of the actual receiver.
@@
These simulations consider both thermal conduction and radiation inside the region of interest. The aerogel filters are treated as uniform, perfect diffusers that transmit, radiate, and reflect uniformly in all directions. Simulations are performed iteratively for different emissivity values ranging from 0 to 0.3. Figure 12 shows the temperature and the total emitted power of the bottom aerogel filter at the 4 K stage as a function of emissivity. The blue line and the orange line represent the result in a three-filter configuration and five-filter configuration respectively. The red dashed horizontal line indicates the upper limit of load (0.1 W) required by the CLASS experiment.
@@
Filter configurations Filter Temperature (K) 4 K Baseplate Temperature (K) 4 K Baseplate Load (W)
@@ VII Cryogenic performance of Aerogel filters
Any scattered stray light is absorbed by the blackened inner shells and baffles on both the 60 K and 4 K stages. The spacing and height of the baffles are set such that they are geometrically black, and all shell interiors are blackened with a mixture of Stycast epoxy and carbon lamp black.
@@
where $T$ is the temperature in Kelvin and $P$ is the power in watts. The 4 K baseplate load inferred from the above load curve for each filter configuration is summarized in the rightmost column in Table 3. Note that this inferred power is much higher than the power emitted by the bottom 4 K filter as simulated by COMSOL. This could be because the thermal load from the window reached the baseplates through other paths, such as the gap between shields of adjacent stages, instead of going through the aerogel filter only. Another possible reason is that there are tiny air bubbles present in the aerogel filters that make them slightly more transmissive in bulk than optical measurements suggest, allowing more load to pass onto the baseplates than expected.